OpenStudy (anonymous):

HELP! how to solve system of equations x-2y+z=-1 y+2z=-4 x+y+3z=-5 what do we do when there is a missing variable??? :(

OpenStudy (anonymous):

what are you solving for??

OpenStudy (anonymous):

x,y, and z

OpenStudy (anonymous):

I'm not sure what to do since the x is missing.

Directrix (directrix):

Write the system as: x - 2y +z =-1 0x +y+ 2z=-4 x + y + 3z = -5 In the first and third equations, eliminate the variable x. Then, you will have a system of 2 by 2 equations in the unknowns y and z. Of those two equations, eliminate a variable and solve for the other. If you are allowed to use technology, take a look at the online 3 x 3 linear equation system solver at: http://math.bd.psu.edu/~jpp4/finitemath/3x3solver.html @mathcalculus

Directrix (directrix):

Question?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

im trying to solve it step by step

Directrix (directrix):

x - 2y +z =-1 0x +y+ 2z=-4 x + y + 3z = -5 x - 2y +z =-1 x - 2y +z =-1 x + y + 3z = -5 --multiply through by -1 -> -x - y - 3z = 5 ---------------- -3y - 2z = 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got a little lost. now i know why

OpenStudy (anonymous):

i got that, so what about the second equation?

Directrix (directrix):

@mathcalculus Your Task: Solve this system - Just add them. -3y - 2z = 4 y+ 2z=-4 -----------------

OpenStudy (anonymous):

it bothers me that the 0x is there.

Directrix (directrix):

Don't look at it. When I eliminated the "x" in equations #1 and #3, the 0x was not an issue.

OpenStudy (anonymous):

i got this: -2y=0

OpenStudy (anonymous):

but it can't be. it's undefined.

Directrix (directrix):

Note that here: Solve this system - Just add them. -3y - 2z = 4 y+ 2z=-4 ----------------- I excluded the 0x term. It is just a placeholder of sorts.

OpenStudy (anonymous):

0/2?

OpenStudy (anonymous):

yeah i know.

Directrix (directrix):

-2y = 0 y= 0/2 y = 0--> You have division by 0 and division into 0 mixed up.

OpenStudy (anonymous):

so y is 0?

OpenStudy (anonymous):

okay i got that step right.

OpenStudy (anonymous):

z=2

OpenStudy (anonymous):

x=-6

OpenStudy (anonymous):

it's wrong... :(

Directrix (directrix):

It is wrong. But, unless you show your steps, I cannot help you find the error.

OpenStudy (anonymous):

z=-2

OpenStudy (anonymous):

ok hold on.

OpenStudy (anonymous):

i got it right! x=1 y=0 z=-2 ! =]

OpenStudy (anonymous):

thanks so much! i really aprreciate it!

Directrix (directrix):

You are welcome.